A parallel-plate air-filled capacitor having area 42.0 cm2 and plate spacing 1.10 mm is charged to a potential difference of 570 V. Find (a) the capacitance, (b) the magnitude of the charge on each plate, (c) the stored energy, (d) the electric field between the plates, (e) the energy density between the plates.
Answer #1

Given

Area of parallel plate capaciotr A = 42 cm 2

distance b etween two plates d = 1.10 mm

Potential difference across the plates V = 570 V

1) Capacitance of the capacitor

               C = εo A /d

                 = ( 8.85*10-12 ) ( 42*10-4 ) / ( 1.1*10-3)

                   = 33.79 pF  

2)    Charge on each plate

             Q = C V

                 = (33.79*10-12)(570)

                 = 19.26 nC

3)Energy stored in the capaciotr

         U = ( 1 /2) C V2  

           = (0.5)(19.26*10-12)(570)2     

            = 3. 128 μJ

4)Electric field between the plates

           E = V / d

               = 570 / 1.1 *10-3

                = 5.18*105 V/ m

5)energy density

         U = ( 1/2) εo E 2  

             = ( 0.5) (8.85*10-12 ) (5.18*105 ) 2

              = 1.1873 *103   ( J / m 3 )